OpenStudy (anonymous):

how do i factor: f(x) = 3x^3 - 12x^2 - 15x

OpenStudy (whpalmer4):

Well, first I would simplify the polynomial by noting that all three coefficients are multiples of 3, so divide them all by 3, giving \[f(x) = 3(x^3-4x^2-5x)\]There's also an x in each term that can be factored out\[f(x)=3x(x^2-4x-5)\] Now to factor \[x^2-4x-5\]Remember that \[(x+a)(x+b) = x^2+bx + ax + ab = x^2 + (a+b)x + ab\]We've got that same form, if we let (a+b) = -4 and ab = -5. Figure out two numbers that fit that constraint, and you're done!

OpenStudy (whpalmer4):

Our final factoring would be \[f(x) = 3x(x+a)(x+b)\]

OpenStudy (anonymous):

so would 3x(x-4) (x+1) work?

OpenStudy (whpalmer4):

Close! What is (x-4)(x+1)? x^2-4x + x - 4 = x^2 -3x -4 Hint: -4*1 = -4, not -5 :-)

OpenStudy (anonymous):

so its x(-5) (x+1) right?

OpenStudy (whpalmer4):

Yep! Full factorization is \[f(x)=3x(x-5)(x+1)\] and if f(x) = 0, it will have solutions at x = 0, x = 5, x = -1.

OpenStudy (anonymous):

yep thanks so much!